1. Suppose the slope of line AB is zero. What is the slope of Line CD given that
a. Line CD is parallel to AB?
b. Line CD is perpendicular to AB?
2. Suppose the slope of Line LT is undefined. What is the slope of Line RD given that
a. Line RD is parallel to line LT?
b. Line RD is perpendicula to Line LT?
3. Tell whether each statement is always, sometimes , or never true.
a. The slope of the line with positive x- and y-intercepts is positive.
b. The slope of the line with negative x- and y-intercepts is negative.
4. Slopes are used in the planning and construction of the roof of a building. The slope of the roof is called pitch. A carpenter must know the pitch of a roof to cut rafters of the correct lenght.
A. The rise of the roof is 14 ft. and the span is 28 ft. Find the pitch.
B. The pitch of the roof Mang Itoy is building if 5 to 12. Half of the spam id 16 ft. What is the rise?
C. About how long should a rafter for the roof decribed in letter A be?

Answer by stanbon(75887)

1. Suppose the slope of line AB is zero.
What is the slope of Line CD given that
a. Line CD is parallel to AB? : zero
--------------
b. Line CD is perpendicular to AB?: undefined
--------------
2. Suppose the slope of Line LT is undefined. What is the slope of Line RD given that
a. Line RD is parallel to line LT?: undefined
--------------
b. Line RD is perpendicula to Line LT?: zero
--------------
3. Tell whether each statement is always, sometimes , or never true.
a. The slope of the line with positive x- and y-intercepts is positive.: never
--------------
b. The slope of the line with negative x- and y-intercepts is negative.: always
--------------
4. Slopes are used in the planning and construction of the roof of a building. The slope of the roof is called pitch. A carpenter must know the pitch of a roof to cut rafters of the correct length.
A. The rise of the roof is 14 ft. and the span is 28 ft. Find the pitch.:14/14
I am assuming the span is from one side of the roof to the other.
---------------------
B. The pitch of the roof Mang Itoy is building is 5 to 12. Half of the span is 16 ft. What is the rise? : 5/12 = x/16 ; x = 20/3 = 6 2/3
----------------------
C. About how long should a rafter for the roof decribed in letter A be?
length = sqrt(14^2+14^2) = 14sqrt(2) = 19.798 ft.
=======================
Cheers,
Stan H.